Helen has been working at the same company in Tampa since she graduated from college. The company has provided Helen with many opportunities, but Helen worries that her salary of $100,000 has not kept pace with what she could earn if she were to change companies. Helen is ready to make a move, and she has been researching the job market. Helen is considering a transfer with her current employer to a division in Atlanta to be closer to her aging parents and her siblings Helen looked up the city indices to help her with the financial implications of her next move. The city indices are: San Francisco: 135 Tampa: Atlanta: 126 103 Consider the following scenarios and answer the questions that follow. Do not round intermediate calculations. Round your answers to the nearest dollar. San Francisco Atlanta Helen found a position with a competitor in San Francisco that is very similar to what she does in Tampa. The salary range is $91,000 to $101,000 How does this compare to Helens current salary? Is Helens current employer paying her a competitive wage? Helen is considering applying for a position that her current employer has posted in a division in Atlanta. The salary range is similar to what she is making now but she knows that the standard of living is lower in Atlanta than it is in Tampa so her current salary of $100,000 should actually have more purchasing power in Atlanta If Helen earns a salary of $100,000 while living in Atlanta, she will be able to purchase an amount of goods and services equal to what amount in Tampa? The middle of the competitors salary range would be equivalent to in Tampa
Answer #1

Middle of Salary in San Francisco = ($91,000+$101,000)/2 = $96,000

Equivalent Salary in Tampa = City Index of Tampa*Salary in San Francisco/City Index of San Francisco

= 126*96,000/135

= $89,600

Since the current employer is paying $100,000, the current employer is paying more than the competitor

Goods and Services in Atlanta, equivalent amount in Tampa = 126*100,000/103

= $122,330.10